### v2

# A Systematic Approach to Predict the Impact of Cybersecurity Vulnerabilities Using LLMs

Abstract
Vulnerability databases, such as the National Vulnerability Database (NVD), offer detailed descriptions of Common Vulnerabilities and Exposures (CVEs), but often lack information on their real-world impact, such as the tactics, techniques, and procedures (TTPs) that adversaries may use to exploit the vulnerability. However, manually linking CVEs to their corresponding TTPs is a challenging and time-consuming task, and the high volume of new vulnerabilities published annually makes automated support desirable.
This paper introduces Triage, a two-pronged automated approach that uses Large Language Models (LLMs) to map CVEs to relevant techniques from the Att&ck knowledge base. We first prompt an LLM with instructions based on MITRE’s CVE Mapping Methodology to predict an initial list of techniques. This list is then combined with the results from a second LLM-based module that uses in-context learning to map a CVE to relevant techniques. This hybrid approach strategically combines rule-based reasoning with data-driven inference. Our evaluation reveals that in-context learning outperforms the individual mapping methods, and the hybrid approach improves recall of exploitation techniques. We also find that GPT-4o-mini performs better than Llama3.3-70B on this task. Overall, our results show that LLMs can be used to automatically predict the impact of cybersecurity vulnerabilities and Triage makes the process of mapping CVEs to Att&ck more efficient.
I Introduction
Vulnerability databases, such as the National Vulnerability Database (NVD),111 https://nvd.nist.gov help cybersecurity professionals keep track of publicly disclosed vulnerabilities (CVEs),222 https://cve.org along with their corresponding labels. These labels include vulnerability descriptions, severity scores (CVSS), weakness classifications (CWEs),333 https://cwe.mitre.org and affected products, which help security experts categorize vulnerabilities relevant to their own systems. A comprehensive overview of vulnerabilities and their impact enables prioritization of defensive measures that need to be taken [1]. To protect organizations from harmful attacks and assess their security posture, security analysts routinely monitor the NVD for vulnerabilities that are added or updated.
While vulnerability databases like NVD provide many technical details, they do not provide any information on how these vulnerabilities may be exploited in real-world attacks. The Att&ck knowledge base444 https://attack.mitre.org aims to address this gap by documenting the tactics, techniques, and procedures (TTPs) of adversaries that have been observed in reported attacks [2]. Understanding how an attacker operates helps formulate a strategy for addressing existing vulnerabilities. Unfortunately, the NVD does not currently label vulnerabilities with their corresponding TTPs, which means that there is no direct connection between the NVD and the Att&ck knowledge base. As a consequence, assessing the risks associated with a given vulnerability and identifying the most relevant TTPs remains a manual, expertise-driven task – making the process both time-consuming and prone to various errors and oversights.
To mitigate this challenge, [3] created the CVE Mapping Methodology (CMM) [3] which guides the mapping of CVEs to their relevant Att&ck techniques or tactics. However, while the CMM is inherently valuable in assisting security analysts with manual mapping, the sheer volume of vulnerabilities published each year calls for the development of a more automated solution to keep CVEs aligned with their corresponding Att&ck tactics and techniques.
We present a novel approach that combines the CMM with Large Language Models (LLMs) to automatically map CVEs to Att&ck techniques. LLMs have demonstrated strong in-context learning capabilities, enabling them to perform complex tasks with only a few task-specific training examples provided as part of their prompt [4]. This makes in-context learning particularly well-suited for cybersecurity applications, where labeled data are scarce and costly to produce.
Contributions: The main contribution of this paper is Triage (“Technique Ranking via In-context learning And Guided Extraction”), a two-pronged approach that automatically maps CVEs to their corresponding Att&ck techniques. We first prompt an LLM using guidance from the CMM to generate a list of relevant Att&ck techniques. Next, we instruct the LLM to use in-context learning on existing CVEs and their mapping to Att&ck techniques to generate a second list of relevant techniques. Finally, we combine the information from both lists into an overall prediction. This hybrid approach strategically combines rule-based mapping from the CMM with data-driven in-context learning based on already labeled CVEs. We empirically evaluate Triage using a MITRE dataset that maps existing CVEs to corresponding Att&ck techniques [5], and we compare our approach to the current state-of-the-art. Moreover, we assess performance across two LLMs, the proprietary GPT-4o-mini and the open-weight Llama3.3-70B, and conduct an ablation study to analyze the contributions of various parts of the prompt used for in-context learning. Our evaluation shows that Triage outperforms the state-of-the-art, and the proprietary model GPT-4o-mini outperforms the open-weight model Llama3.3-70B on this task. We conclude that LLMs can be used to automatically predict the impact of cybersecurity vulnerabilities and Triage makes the process of mapping CVEs to Att&ck more efficient. We include the various prompts in the Appendix and provide a replication package [6].
II Background & Related Work
II-A CVE Mapping Methodology (CMM)
The CMM defines an approach to map CVEs to Att&ck techniques [3]. These mappings help determine the impact of a vulnerability and how it can be exploited in the real world. The CMM distinguishes attack techniques related to three attack phases: First, a vulnerability enables the attacker to use a specific exploitation technique. Second, the exploit grants the attacker certain capabilities, known as the primary impact. Third, the attacker can gain further capabilities, leading to the secondary impact. These phases, illustrated in Figure 1, are adapted from the intrusion kill chain introduced by [7] [7]. In each phase, the defender can detect and prevent the attack using various strategies.
A vulnerability can be mapped to multiple Att&ck techniques, and a mapping is a one-to-one relationship between a vulnerability and an attack technique, labeled with one of the three possible mapping types (exploitation technique, primary or secondary impact). The CMM provides five methods to analyze a CVE. A single method will not necessarily provide all relevant attack techniques across the three mapping types, but the methods can be combined to obtain more complete mappings. The five methods are as follows:
II-A-1.
Vulnerability type: This method supports mapping a CVE based on its vulnerability type. It defines a table with relevant attack techniques (or instructions on how to find them) for 27 of the most common vulnerability types. These vulnerability types are based on CWEs; however, some of them group multiple CWEs to keep the number of alternatives manageable. An example is Cleartext Transmission of Sensitive Information which is mapped to T1040: Network Sniffing as exploitation technique, T1552: Unsecured Credentials as primary impact, and T1078: Valid Accounts as secondary impact.
II-A-2.
Functionality: This method defines a table that classifies a CVE by the functionality a successful exploit provides. For example, the ability to Delete Files, which maps to the primary impact T1485: Data Destruction, and secondary impact T1499: Endpoint Denial of Service.
II-A-3.
Exploitation Technique: This method asks the question: “what steps are needed to exploit this CVE?”, and defines a set of possible answers that are linked to relevant attack techniques. For example, if the user needs to visit a malicious website, the exploitation technique can be mapped to T1189: Drive-by Compromise.
II-A-4.
Affected Object: This method defines a table of objects that may be affected by an exploit, and provides mappings to associated attack techniques. Objects include hardware, firmware, operating systems, applications, services, etc. For example, Network-based Application is mapped to exploitation techniques T1040 Network Sniffing, and T1059: Command and Scripting Interpreter.
II-A-5.
Tactic: This method defines a table that associates Att&ck tactics with generic exploitation techniques. Using this method is only recommended when other methods do not provide a relevant answer. For example, the tactic Lateral Movement is associated with the technique T1210 Exploitation of Remote Services.
II-B Mining ATT&CK Techniques From Linked Data
Cybersecurity data sources, such as CVE, CWE, CAPEC and Att&ck, capture related and at times overlapping concepts. These data can be mined to uncover mappings from CVEs to attack techniques. [8] [8] investigated this route by creating BRON, a graph database that combines these data sources and unifies the relations between them. The graph can be queried, and edges can be traversed as bi-directional links to identify patterns, discover new connections, and gain deeper insights into threats and vulnerabilities. The authors find that the links are often sparse; for example, not every CVE or CWE is linked to an associated attack technique. In a follow-up paper [9], the authors extended BRON with additional sources and propose an approach to predict missing links. However, the focus was on predicting links to neighboring nodes (e.g CWE to CAPEC), and they do not predict multi-step links, such as from CVE to Att&ck techniques.
A clear benefit of establishing such linked data is the reduced cost of labeling. However, in addition to the sparsity challenge, linking through CWEs – which are more generic than CVEs – may lead to a loss of context, for example, regarding attack steps indicated by the CVE description. As a consequence, the techniques mined through such links will cover much more than what would be mapped through the CMM. This aspect is investigated by [10] [10], who unify data sources similar to BRON, but enforce stricter conditions on the CWEs, showing that relations between multiple CWEs may lead to an exploding attack surface. In one example, they find that BRON maps to 84 different techniques, whereas their more strict approach maps to only two. They also opt for uni-directional linking, which they find has considerable advantages on analysis speed.
II-C Automating the Mapping from CVE to ATT&CK
Data-driven methods: Att&ck distinguishes various levels, and the literature includes approaches for automating the mapping from CVE to Att&ck tactics [11, 12], Att&ck techniques [13, 14], and both [15]. Tactics are described at a higher level than techniques, for example, Initial Access describes the adversary goal of getting into a network. How the adversary enters the network is then described through attack techniques associated with the tactic, for example T1566: Phishing. Our review will focus on automated mapping from CVEs to techniques, which is the goal of Triage.
[13] [13] compare multi-labeled supervised approaches with and without data augmentation. They include classical machine learning (Support Vector Machines and Naive Bayes Classifiers), deep learning models trained from scratch, and fine-tuning of pre-trained models based on variants of BERT architectures. The classical models are used to establish a baseline and help interpret relevant features. Not surprisingly, these models do not perform at the same level as the deep learning models, particularly the pre-trained BERT-based architectures. The authors perform data augmentation using the TextAttack555 https://textattack.readthedocs.io/en/master framework to generate adversarial examples by transforming the input text. They find that data augmentation boosts the performance of the deep learning models, but not the classical models. Their two best-performing approaches relied on SciBERT and SecBERT to encode the representations of the CVE text descriptions.
[14] [14] propose SMET, which uses Semantic Role Labeling (SRL) and sentence embedding to map CVEs to techniques. SMET does not need CVEs labeled with attack techniques and considers all 185 techniques that were in Att&ck at the time of the study. It consists of three steps: (1) attack vector extraction through Semantic Role Labeling (SRL), (2) attack vector representation using Att&ck BERT, and (3) attack technique prediction using Logistic Regression trained on the embeddings produced by Att&ck BERT.
The attack vector extraction of SMET identifies subject-verb-object structures from the input text using an SRL model from AllenNLP. SMET is trained on attack vectors extracted from technique descriptions and procedure examples from Att&ck. An attack vector is extracted if either the words “adversary,” “vulnerability,” or “user” are among the subject words, or the verb is either “allow,” “lead,” or “result.” The authors do not discuss why these words were selected, or to what extent they cover the attack vectors found in CVEs.
The authors develop their own Att&ck BERT model which is fine-tuned using the SIAMESE network to optimize the similarity between two attack vectors. In this context, two attack vectors are considered similar if they either share the same objective or help to achieve the same technique. The authors find that SMET using their own Att&ck BERT outperforms instances of SMET that use SecBERT or SBERT.
[16] [16] also apply SRL to this task. However, in contrast to SMET, which exclusively uses the attack vectors extracted from CVE descriptions, they also consider the surrounding context (the full CVE description) as input to the classification model. Moreover, they incorporate attack functionality prediction as an intermediate step towards predicting the attack techniques. This step partially draws on the CMM, as predicted functionalities can be mapped to ATT&CK techniques using the methodology’s predefined associations (method II-A-2). However, functionality mapping represents only one of the five methods outlined in the CMM, and it does not provide comprehensive coverage of all techniques. In contrast, our proposed Triage approach leverages the full range of methods from the CMM, along with labeled examples. Moreover, it is the only automated approach that explicitly models the three mapping types from the CMM (exploitation techniques, primary and secondary impacts), providing additional context to each mapping, and offering more complete insight on how a vulnerability is exploited.
[17] [17] propose an approach that first maps a CVE to a CWE, and then uses both the CMM and linked data (Section II-B) to map from CWE to attack techniques. Their approach, which is based on [18] [18], performs feature extraction on CVE descriptions experimenting with both word embeddings and TF-IDF. Using the extracted features, they train a set of neural networks to predict CWEs from CVEs. In addition to approach described above, the authors experiment with associating CVEs to techniques directly using a similarity search approach with Doc2Vec [19], however this alternative approach performs significantly worse.
LLMs: As part of evaluating SMET, [14] [14] study whether “one of the most popular chatbots” can map CVEs to techniques. They provide this chatbot (left unnamed in the paper) with a CVE description and ask to map it to Att&ck techniques. They also apply the chatbot as the extraction component in SMET. In both cases, SMET consistently outperforms the LLM. The authors experimented with three different prompts, all of them relatively short, limited to the CVE description, and two sentences for instructions.
[20] [20] investigate the use of LLMs for two tasks: (1) mapping from CVE to CWE, and (2) mapping from CVE to Att&ck. The authors experiment with weak prompts, which are straightforward questions related to the task, and strong prompts, which use a chain-of-thought process where they first ask five options for possible answers, and then ask for the final answer based on these. The authors consider the performance of the LLM in all experiments to be unsatisfactory, which they attribute to poor quality of the data, as well as the LLM lacking a conceptual understanding of Att&ck.
[15] [15] experiment with different fine-tuning approaches to make LLMs more domain-specific. Their approach, VTT-LLM, shows a clear improvement compared to [20] [20]. The first stage of their two-stage approach uses data from BRON to prompt a generative LLM aiming to capture the links from CVE via CWE and CAPEC to Att&ck techniques and tactics and turn a CVE description into a technical attack description. Their best approach uses a template chain, explicitly capturing these links by asking the LLM to “think step by step” over each data point. In the second stage, they use embeddings to associatively match the generated attack description with technique descriptions derived from the Att&ck knowledge base.
Finally, [21] [21] explore an idea related to the present paper, describing how they prompt ChatGPT based on the CMM and few-shot examples. However, their contributions are difficult to evaluate, as their prompts only define overall mapping goals as the context and include a CVE description in the user prompt, omitting details for the five methods of the CMM, as well as the few-shot examples of labeled CVEs. The authors mention consolidating the primary and secondary impacts due to the difficulty of establishing the relationship and ordering between them.
Comparative Analysis and Reproduction Studies: [22] re-implement and compare five related studies [22]. In addition, the authors add hyperparameter tuning and data augmentation to each approach. They provide an overview of the differences in results but do not compare in more detail. As such, two challenges remain: First, the evaluations often employ different metrics that cannot be directly compared. Second, there is significant variation in the number of attack techniques covered by the different datasets, varying from 17 to 52 in the considered literature. Training and testing the approaches from each study on standard datasets using the same metrics would be interesting for a more detailed comparison. Finally, although the authors report strong results for their data augmentation, closer analysis of their procedure shows that training samples can leak into their test set because the data are augmented before splitting out the test set.
III The Hybrid Triage Approach
Our goal is to take a CVE as input and automatically map it to relevant attack techniques defined in Att&ck. We propose a hybrid approach with multiple separate LLM invocations to capture the various aspects of each CVE. Finally, we combine the output of individual runs to obtain a list of technique predictions per mapping type. Triage is illustrated in Figure 2 and consists of two main components, the Methodology Mappers, which implement rule-based mapping and prompting following the CMM, and the In-Context Learner, which implements many-shot in-context learning based on already labeled CVEs.
III-A Methodology Mappers
As input, the model takes a CVE with its corresponding description. Using a template-based approach, we create prompts for each of the five methods of the CMM, and run these separately through the LLM. The prompt templates used for each method are detailed in Appendices -A to -E.
Vulnerability type: Vulnerability types are closely related to CWEs. Thus, we can match vulnerability types with corresponding CWE types, and use the CWE description as a description of the vulnerability type. Based on this list of descriptions, we ask the LLM to identify which vulnerability type matches the CVE description the best. Although occasionally multiple CWEs can be present in a CVE record, the common practice is to map a CVE to a single CWE. For this reason, we instruct the LLM to provide only one vulnerability type. Predicted vulnerability types are then mapped to attack techniques using the CMM table for this method (II-A-1).
Functionality: Our prompt first explains what a functionality is, and lists possible functionality types in the CMM. Descriptions of functionality types are not included in the CMM, but we reuse the descriptions provided by [16] [16]. We allow the LLM to predict multiple functionalities, which are then mapped to attack techniques using the CMM table for this method (II-A-2).
Exploitation Technique: The CMM defines eight top-level questions for this method, which in two cases are followed by a multiple-choice sub-question. First, we provide the LLM with the top-level questions as “Yes/No” questions. When relevant, we provide it with the sub-questions and answers to choose from. The model can answer “Yes” to multiple top-level questions. Aggregated answers are mapped to attack techniques using the CMM rules for this method (II-A-3).
Affected Object: The prompt includes a description of affected objects along with possible categories. Descriptions of individual object types are not provided in the CMM, so we manually created a list of descriptions for the various object types in this method. We allow the model to predict multiple affected objects which are then mapped to attack techniques using the CMM table for this method (II-A-4).
Tactic: The prompt contains a general definition of a tactic according to the Att&ck knowledge base, and provides brief descriptions of each possible type. We allow the model to predict multiple tactics which are mapped to attack techniques following the CMM table for this method (II-A-5).
III-B In-Context Learner
The In-Context Learner creates attack mappings based on relevant features and in-context examples without any direct reference to the CMM. The prompt was developed iteratively, adding relevant features and refining these based on the response. The prompt template is detailed in Section -F. It begins with categorizing an attack in three steps: Exploitation technique, Primary Impact and Secondary Impact. Next, we provide all enterprise attack techniques as classes to choose from. For each technique, we include the attack names and the full descriptions from Att&ck. Then, for each target CVE, we include CVSS scores and CWE types in addition to the CVE description. Next, we include in-context examples from our training set. Each in-context example includes CVE ID, CVE description, attack technique ID, attack name, and mapping type. The last part of the prompt summarizes the task and the output format, in which the LLM is instructed to predict a ranked list of the ten most relevant attack techniques for a given mapping type.
III-C Combining Results Into Final Predictions
To obtain the final attack technique predictions we combine the results from the Methodology Mappers and the In-Context Learner. First, from the Methodology Mappers, the intermediate predictions of each method are aggregated in a list. Next, for each CVE and mapping type, we combine the list of top ten predictions from the In-Context Learner with the aggregated list of Methodology Mappers predictions.
The In-Context Learner is a purely data-driven approach, while the Methodology Mappers mimic a rule-based expert system. By combining these two approaches, we aim to get the best of both worlds in our final predictions.
III-D Research Questions
RQ1:
To what extent can systematically applying the CMM automate the mapping from CVE to Att&ck?
RQ2:
Is there a performance difference between prompting based on the CMM vs. using in-context learning with ground truth labels based on the methodology?
RQ3:
How can we combine the two approaches from RQ2, and to what extent does this lead to performance gains?
RQ4:
What is the practical relevance of Triage in terms of efficacy and quality of the resulting predictions?
IV Experimental Design
IV-A Dataset
This work builds on the KEV (Known Exploited Vulnerabilities) dataset provided by [5] [5]. The dataset contains 296 CVEs labeled with Att&ck techniques, comprising 806 mappings. The distribution of individual techniques over these mappings is visualized in Figure 3. As can be observed, the distribution of mapped techniques is highly imbalanced. Moreover, exploration of the dataset revealed that a significant number of mappings refer to techniques not mentioned in the CMM, and several techniques included in the CMM do not occur in the dataset.666 A total of 14 of 306 exploitation technique mappings, 34 of 256 primary impact mappings, and 58 of 244 secondary impact mappings in the dataset refer to techniques that are not in the CMM. Conversely, from all techniques in the CMM, the dataset uses 11 of 15 exploitation techniques, 16 of 24 primary impacts, and 14 of 18 secondary impacts. Our interpretation is that the annotator must have relied on external knowledge not included in the CMM. The methodology does not prohibit this; however, our Methodology Mapper follows the five CMM methods and will predict corresponding techniques. On the other hand, the In-Context Learner should be able to handle such cases, so we keep these mappings in the dataset.
Table I shows how frequent each mapping type is across the CVEs in the dataset. The table highlights that exploitation technique and primary impact most commonly have one technique while secondary impacts mostly have none which occurs when a mapping type is missing for a CVE.
From the dataset, we extract CVE IDs, attack IDs, attack names, and mapping types. Moreover, we enrich the data by adding CVE descriptions, CVSS scores, CWE name, and CWE description to each CVE. Similarly, for each attack technique, we add Att&ck descriptions. The 806 mappings in the data correspond to 296 CVEs. We split the labeled data in train and test splits with an 80/20 split over the CVEs.
| mapping type | none | one | two | three | three |
| exploitation | 12 | 264 | 19 | 0 | 1 |
| primary | 57 | 225 | 11 | 3 | 0 |
| secondary | 180 | 69 | 21 | 7 | 19 |
IV-B Evaluation Metrics
In addition to evaluating our final predictions, we also evaluate each method (prompts) of the Methodology Mappers along with various configurations of the In-Context Learner. Here we describe the metrics used in these evaluations.
Unranked Classification: The Methodology Mappers generate unranked predictions of how a CVE can be classified into techniques. These are multi-label predictions that cover the three mapping types. We also consider a variation on our In-Context Learner that generates a similar unranked multi-label classification. To score these results, we use the standard evaluation metrics for unranked sets: precision (P), recall (R), and their harmonic mean F1 [23, Ch. 8.3].
We adapt the metrics to multi-label classification by micro-averaging, which places equal weight on individual instances across all classes. We focus on micro-averaging since our data is imbalanced, and we want to avoid that rare classes have a disproportionate influence on the overall metric value. A challenge with using micro-averaging on the highly imbalanced data with many classes is that a model that always predicts the majority class can achieve good performance. To mitigate this, we also analyze performance over classwise predictions, which are computed by framing the multi-label task as multiple binary classification problems, treating each target technique as positive and all others as negative.
Ranked Results: Both the In-Context Learner and the final predictions produce ranked lists of techniques for each of the three mapping types. To score these results, we use evaluation metrics that consider not only whether a technique is correct but also its position in the list [23, Ch. 8.4].
Assume that, for one of the three mapping types, a given CVE is mapped to a (possibly empty) list of associated techniques in the ground truth (the labeled dataset). Let us also assume the model outputs an ordered list of techniques predicted to be associated with the CVE for that mapping type. We can then specify a cut-off threshold and define , abbreviated to P@, as the number of correctly predicted techniques in the top- divided by :
where are the top- techniques from the ones predicted by the model. Similarly, or R@ is defined as:
For Average Precision, or AP, we compute for all possible values of :
where , and denotes the indicator function.
Finally, we compute the Mean Average Precision (or MAP) as the mean of all individual AP scores across all labeled examples in the ground truth dataset, denoted as Q:
The MAP score measures both whether or not a prediction is correct but in addition also considers quality of the ranking.
IV-C Technical Details
We describe here some implementation choices that were not discussed when describing the overall approach in Section III.
Lifting Sub-Techniques: In the In-Context Learner, we lift all sub-techniques to their top-level parent technique to reduce the number of classes that the model must handle and shorten the list of examples in the prompt.
Combining Lists of Ranked Predictions: We use the following strategy for aggregating predictions from the Methodology Mappers and In-Context Learner: (1) we want to avoid overwriting high-ranked predictions from the In-Context Learner with lower-ranked predictions from the Methodology Mappers; (2) we want to keep “None” as prediction from the In-Context Learner when it appears. Thus, we merge Methodology Mapper-generated predictions into In-Context Learner ones in descending order, starting from the last position. We overwrite any lower-ranked In-Context Learner prediction with novel predictions from the Methodology Mappers, except when there is a “None” value at the corresponding position. This way, we get the best from the In-Context Learner, augmented by novel additions from the Methodology Mappers.
IV-D Ablation Study
To investigate how different components of the In-Context-Learner prompt influence the performance, we run a set of ablations over 100 randomly selected CVEs, and focus on the following four features: (1) Attack descriptions provide extra context to the options the model should select from; (2) CVSS metrics include information on how the vulnerability may be exploited and what aspects of the system are affected; (3) CWE types are closely related to vulnerability types and may indicate attack techniques, cf. the CMM (II-A-1); (4) In-context examples enable in-context learning from existing mappings. However, they are generally expensive to create and increase the context window.
IV-E Enabling a Comparison between TRIAGE and SMET
We chose to compare our results to SMET [14], which also produces a ranked list of techniques. This requires a few additional steps. First, we need to collect SMET results for the same dataset as we use. To collect the results, we can reuse the models previously trained by SMET’s authors, as these depend on Att&ck, not on the evaluation dataset.
The next step comes from the fact that SMET does not distinguish between the three mapping types from the CMM. To compare Triage, which does distinguish these mapping types, we perform post-processing of our results to merge the three mapping types into a single category. Since these are ranked lists, the merging is done as follows: The top-ranked technique is “popped” from the list of exploitation techniques and added to the results; then we do the same for the top of the primary impact list, and then the top of the secondary impact list. We repeat this process until all the lists are empty.
V Results and Discussion
| mapping type | mapping method | train | test | ||||
| F1 | P | R | F1 | P | R | ||
| exploitation | vulnerability type | ||||||
| exploitation technique | |||||||
| affected object types | |||||||
| tactic | |||||||
| primary | vulnerability type | ||||||
| functionality | |||||||
| secondary | vulnerability type | ||||||
| functionality | |||||||
| exploitation | MM combined | ||||||
| primary | MM combined | ||||||
| secondary | MM combined |
| exploitation | In-Context Learner | ||||||
| primary | In-Context Learner | ||||||
| secondary | In-Context Learner |
V-A Assessing the Individual Components of TRIAGE
We first explore how well individual mapping methods compare, considering all Methodology Mappers, and the In-Context Learner (ignoring the ranking for now). The upper part of Table II shows the results for individual Methodology Mappers (observe that not all are relevant for each mapping type); the middle part shows them combined, taking the union of all answers per mapping type; and the lower part shows the In-Context Learner. Focusing on the F1 scores, we see that the In-Context Learner consistently outperforms the other methods. The vulnerability type method is far better at predicting primary impact than exploitation technique or secondary impact. As expected, the exploitation technique method is the best among Methodology Mappers for predicting exploitation techniques. The functionality method performs poorly for both primary and secondary impacts on both data splits. (and does not predict exploitation techniques cf. the CMM). The combination of all Methodology Mappers performs well in recall, but poorly in precision. This is in part due to the noise introduced by the poorly performing methods (precision increases when they are removed). An underlying cause here is that a considerable number of techniques mentioned in the CMM are not used in the dataset, so every prediction of these creates a false positive (see also Section IV-A).
V-B Comparing Two Different LLMs for Powering TRIAGE
Moving towards our hybrid approach, we change from unranked to ranked predictions. As a baseline, we will first assess the performance of the In-Context Learner by itself. Next, we will assess the performance of our hybrid combination of the Methodology Mappers with the In-Context Learner. Considering the poor performance of various Methodology Mappers in the previous section, only the following are included in our hybrid: (1) vulnerability type, (2) exploitation technique, (3) affected object. We will also use this opportunity to compare the performance of two LLMs, the proprietary GPT-4o-mini and the open-weight Llama3.3-70B.
The upper part of Table III shows the performance of both LLMs using the In-Context Learner alone. Considering the MAP scores, GPT-4o-mini consistently outperforms Llama3.3-70B across all mapping types and data splits. Looking at recall, GPT-4o-mini always performs better than Llama3.3-70B, except Recall@5 for exploitation technique and Recall@10 for primary impact, where they score equally.
The lower part of Table III shows the performance of Triage. Taking into account the MAP scores, GPT-4o-mini again outperforms Llama3.3-70B in all mapping types and data splits. We see that the performance related to Recall@10 is almost the same between the two, indicating that the hybrid approach evens out the initial differences related to exploitation technique. This suggests that Methodology Mappers extract relevant exploitation techniques that are not identified by the In-Context Learner. For primary or secondary impact, we do not see a similar effect. Moreover, for secondary impact, we see a large difference when comparing Recall@10 with Recall@5 in Table III, which means that many of the extracted techniques are not among the top 5 ranked predictions. This is also reflected in the low MAP scores for secondary impact, indicating that the model is poor at ranking the correct techniques. We also note that Recall@10 increases significantly for secondary impact when comparing the training and test sets for both models. This could be related to the fact that the test set has far fewer unique attack techniques (18) compared to the training set (71), when considering the secondary impacts. Rare techniques are typically more difficult to predict, which we will examine in more detail in the next section. At a high level, the results from Table III indicate that exploitation techniques are the easiest types to predict for the model, followed by primary impacts and finally secondary impacts.
mapping type mapping method train test MAP R@10 R@5 MAP R@10 R@5 exploitation ICL (GPT-4o-mini) ICL (Llama3.3-70B) primary ICL (GPT-4o-mini) ICL (Llama3.3-70B) secondary ICL (GPT-4o-mini) ICL (Llama3.3-70B)
exploitation Triage (GPT-4o-mini) Triage (Llama3.3-70B) primary Triage (GPT-4o-mini) Triage (Llama3.3-70B) secondary Triage (GPT-4o-mini) Triage (Llama3.3-70B)
V-C Ablation Study and Analysis of Classwise Predictions
In addition to Section V-A, which assesses the contributions of Triage components, we conduct an ablation study on the In-Context Learner, gradually removing features from the prompt. As shown in Table IV, the CVSS and CWE features have a limited impact on the various mapping types. For primary impact, including CVSS even reduces recall. Omitting attack descriptions (A) decreases all three metrics for primary and secondary impact. For exploitation technique, omitting attack descriptions increases the MAP by one percentage point, while Recall@5 and Recall@10 decrease. A clear pattern is that exploitation techniques mapping benefits from more in-context examples. Reducing from 235 to 30 examples reduces performance by 7 to 10 percentage points across the metrics. An even larger performance drop is seen when removing all examples. The secondary impact is less influenced by removing features, which is likely due to the generally lower performance for this mapping type.
| mapping type | features | #demo | MAP | R@10 | R@5 |
| exploitation | a+cvss+cwe | 235 | |||
| a+cvss | 235 | ||||
| a | 235 | ||||
| 235 | |||||
| 30 | |||||
| 0 | |||||
| primary | a+cvss+cwe | 235 | |||
| a+cvss | 235 | ||||
| a | 235 | ||||
| 235 | |||||
| 30 | |||||
| 0 | |||||
| secondary | a+cvss+cwe | 235 | |||
| a+cvss | 235 | ||||
| a | 235 | ||||
| 235 | |||||
| 30 | |||||
| 0 |
The heatmap in Figure 4 zooms in on the classwise performance for predicting individual Att&ck techniques. Note that the figure only includes techniques that have at least one correct prediction among any of the mapping types. The chart shows large variations in how well individual techniques are predicted across the mapping types and data splits, ranging from perfect scores to no correct predictions.
V-D Performance comparison of SMET and TRIAGE
When comparing our approach with SMET, we consider CVE descriptions mapped to techniques, without mapping types. Moreover, we have seen in our earlier results that secondary impact techniques are more difficult to predict than primary impact and exploitation technique. Therefore, we explore two scenarios which, respectively, include and exclude those secondary impacts from the ground truth and results.
In Table V, we compare SMET with Triage for the two scenarios. We see that our hybrid approach outperforms SMET across all metrics for both configurations. We also see in Table V, that performance for both approaches improves when secondary impacts are excluded.
| mapping method | secondary impacts | train | test | ||||
| MAP | R@10 | R@5 | MAP | R@10 | R@5 | ||
| SMET | included | ||||||
| Triage* | included | ||||||
| SMET | excluded | ||||||
| Triage* | excluded | ||||||
V-E Qualitative Analysis
For our qualitative analysis, we randomly selected eight CVEs, and found that the predicted mappings were correct for five CVEs, while the remaining three had one or more mispredictions. In addition, we observed that the ground truth for five CVEs had empty values for the secondary impact, and two CVEs had empty values for both the primary and the secondary impact. Due to space limitations, we present four of eight analyzed CVEs in this paper, which were selected as follows: First, we include the three CVEs with mispredictions to investigate how our approach missed the target. Analysis of these CVEs is presented in Table VI, VII, and VIII. Then, to explore the empty values in the ground truth in more detail, we include one CVE with empty values for both primary and secondary impact (Table IX).
Each analysis follows the same pattern: We start by analyzing the differences between the ground truth and the predictions of Triage. Next, we consider to what extent the ground truth labels follow common patterns in the data. In the end, we manually re-evaluate the CVE description using CMM to better understand potential differences between our predictions and the ground truth.
Looking at Tables VI and VII, we see that the first one has no secondary impact, even when the description directly mentions a vulnerability type with a secondary impact. The second one has as much as twelve secondary impacts, but these are hard to trace back to the CMM, and are sparse in the training data. In Table VIII, we see a case where the ground truth has a primary impact that cannot be derived through the CMM. Also noteworthy is Table IX, where Triage correctly identifies plausible primary and secondary impacts (following CMM rules) that are missing from the ground truth.
Description: Adobe ColdFusion Update 5 and earlier versions, ColdFusion 11 Update 13 and earlier versions have an exploitable Deserialization of Untrusted Data vulnerability. Successful exploitation could lead to arbitrary code execution. Exploitation Technique: T1203: Exploitation for Client Execution Primary Impact: T1190: Exploit Public-Facing Application, T1133: External Remote Services Analysis: The primary impact T1133 is not found by our approach. This may be caused by the fact that T1133 most commonly appears as exploitation technique in the ground truth, and this CVE is the only case with T1133 as primary impact. Our approach includes T1190 as exploitation technique instead of primary impact. T1190 is most commonly used as exploitation technique in the ground truth, but three times it occurs as primary impact. Deserialization of Untrusted Data is mentioned in the CVE description, and is a vulnerability type in the CMM that points to T1059 as primary impact, which is not among the ground truth labels.
Description: Zoho ManageEngine ADSelfService Plus version 6113 and prior is vulnerable to REST API authentication bypass with resultant remote code execution. Exploitation Technique: T1190: Exploit Public-Facing Application Primary Impact: T1505: Server Software Component Secondary Impact: T1573: Encrypted Channel, T1560: Archive Collected Data, T1087: Account Discovery, T1070: Indicator Removal, T1047: Windows Management Instrumentation, T1003: OS Credential Dumping, T1136: Create Account, T1218: System Binary Proxy Execution, T1003: OS Credential Dumping, T1140: Deobfuscate/Decode Files or Information, T1027: Obfuscated Files or Information, T1505: Server Software Component Analysis: Our approach correctly predicts the exploitation technique T1190. It misses T1505 in the primary impact, but includes it in the secondary impact. Interestingly, the ground truth for this CVE includes as many as 12 secondary impact techniques, indicating that there are several different paths open to the attacker. Among these 12, only T1505 is included in our predictions. Some of the secondary impacts are quite rare in the training data, but all of them occur at least twice, which means there is always at least one in-context example with the same technique available. Earlier, we saw that secondary impact predictions already have the lowest performance; a single or just a few in-context examples may not be enough to predict this mapping type. Finally, when manually mapping the CVE description using CMM, we cannot find answers for the primary and secondary impacts.
Description: Microsoft Windows Support Diagnostic Tool (MSDT) Remote Code Execution Vulnerability Exploitation Technique: T1566: Phishing Primary Impact: T1204: User Execution Secondary Impact: T1059: Command and Scripting Interpreter Analysis: Our approach does not predict T1566: Phishing as exploitation technique. The first and the second ranked exploitation technique are instead T1204: User Execution and T1190: Exploit Public-Facing Application. T1566 appears seven times as exploitation technique in the training data so there are at least some in-context examples available. Using the CMM, Operating System among affected object types with T1574: Hijack Execution Flow seems relevant, the attack technique is also included among the predictions. Moving on to the primary impact, T1204 is not included among the primary impact predictions where T1059 is ranked at top. T1204 is used as primary impact two times meaning there is a single in-context example available. Using the CMM, we are not able to map the CVE description to a primary impact. Regarding the secondary impact, our approach predicts the correct technique T1059 as the fifth ranked technique.
Description: Server Side Request Forgery in vRealize Operations Manager API (CVE-2021-21975) prior to 8.4 may allow a malicious actor with network access to the vRealize Operations Manager API can perform a Server Side Request Forgery attack to steal administrative credentials. Exploitation Technique: T1190: Exploit Public-Facing Application Analysis: Our approach correctly identifies T1190 as an exploitation technique. The ground truth has no primary and secondary impacts, which is surprising considering that the term Server Side Request Forgery in the description is a vulnerability type in the CMM pointing to T1133: External Remote Services as exploitation technique, T1090: Proxy as primary impact, and both T1135: Network Share Discovery and T1005: Data from Local System as secondary impacts. Other potential relevant attack techniques predicted by the In-Context Learner include primary impacts T1555: Credentials from Password Stores, T1078: Valid Accounts, and T1068: Exploitation for Privilege Escalation, and secondary impacts T1555, T1078, and T1552: Unsecured Credentials.
V-F Addressing our Research Questions
RQ1: To what extent can systematically applying the CMM
automate the mapping from CVE to ATT&CK?
To address the first research question, we assess the Methodology Mappers in our approach, individually prompting each method of the CMM.
Although there are large variations in performance across mapping types and mapping methods, the results in Table II suggest that using individual components in Methodology Mappers is not sufficient to reach good performance.
The combination of all methods yields better results, but still falls short in performance.
As mentioned when discussing the dataset (Section IV-A), several mappings are related to techniques not present in the methodology, and this is most prevalent for primary and secondary impacts.
RQ2: Is there a performance difference between prompting based on the CMM vs. using in-context learning with ground truth labels based on the methodology?
The answer here is a resounding yes. Table II shows that the In-Context Learner results in higher F1 performance than any of the Methodology Mappers individually or all combined.
Furthermore, the In-Context Learner can provide additional predictions through a ranked approach providing a list of most likely answers.
We consider the In-Context Learner the most impactful component of our approach, given the demonstrated improvement in performance, and the fact that a considerable number of annotations do not strictly follow the CMM.
RQ3: How can we combine the two approaches from RQ2,
and to what extent does this lead to performance gains?
We experimented with two architectures for our hybrid approach:
(1) The first includes the output from Methodology Mappers in the prompt of the In-Context Learner with in-context examples.
(2) The second uses the two-pronged approach from Figure 2, invoking the In-Context Learner for each mapping type.
The first approach “confused” the LLM in the sense that the answer of the Methodology Mappers took precedence, and no other techniques were added based on the in-context examples.
The second approach shows a performance increase for the exploitation technique but not for the primary or the secondary impact (Table III).
RQ4: What is the practical relevance of TRIAGE in terms of efficacy and quality of the resulting predictions?
There are three aspects to answering this question.
First, we compare our work with SMET, the current state-of-the-art in automated mapping.
There is a clear improvement in predictive performance for Triage compared to SMET (Table V).
Moreover, our approach provides additional cybersecurity context, by classifying into CMM mapping types.
Next, our qualitative analysis found that Triage included the correct mappings for five of eight CVEs. We observed that even if the CVE directly mentions a Vulnerability Type present in the CMM, the annotator may have selected other attack techniques. In this case (CVE-2021-21975, Table IX), Triage predicted plausible impacts following the CMM, but these were missing from the ground truth. Our qualitative analysis supports our quantitative analysis that predicting secondary impacts is a difficult task for the LLMs. The lack of data to learn from is a considerable factor here.
The last aspect to consider is what it costs to use Triage to analyze a CVE and predict mappings for all three mapping types, and how long it takes? The combined number of tokens to map a CVE using Triage is in the order of 400k input tokens, divided in three prompts for the Methodology Mappers and three for the In-Context Learner. Sequentially, the total process takes on average 50 seconds for GPT-4o-mini (OpenAI API) and 2m50s for Llama3.3-70B (Lambda AI API) and produces 2k output tokens.777 Parallelization speedups depend on user rate limits; we get around 50%. Thus, the overall Triage mapping costs per CVE currently correspond to US$0.07 for GPT-4o-mini and US$0.05 for Llama3.3-70B.
Comparing again with SMET, run-times or costs are unfortunately not reported by its authors, and since we do not train its models, we cannot compute these costs ourselves. However, when we run inference with SMET on an AMD EPYC 7413 CPU with 2.85 GHz, the processing time averages around 0.6 seconds per CVE. So, purely looking at run times, SMET outperforms Triage, although as we have seen, it does so at a lower predictive performance.
Based on these findings, we argue that Triage has practical relevance, where the Methodology Mappers select specific techniques when there is a clear path in the methodology, and the In-Context Learner provides the flexibility needed to fill in the gaps in the methodology. Moreover, using Triage to support analysts makes the process of mapping CVEs to Att&ck more efficient, at low cost and runtime overhead.
V-G Threats to Validity
Internal validity: Data leakage is a threat and comes in two forms: (1) The LLMs may have seen the mapping of the CVEs. The dataset’s release after the cut-off dates for the LLMs888 Dataset: 2025-03-12; Cut-off dates are resp. Oct 2023 for GPT-4o-mini, released 2024-07-18, and Dec 2023 for Llama3.3-70B, released 2024-12-06. mostly mitigates this, but mapped examples might have been online before. (2) Many CVE descriptions literally include the vulnerability type, e.g., “SQL Injection”, which can lead the In-Context Learner to matching examples. We do not consider this a threat in Triage since the Methodology Mappers follow CMM rules that do the exact same thing.
External validity: One threat is generalizability beyond our ground truth. Performance may differ on non-KEV CVEs, and in other domains. The lack of more labeled data makes it difficult to mitigate this threat. There is also the threat of provider and model dependency. We evaluate two LLMs, the proprietary model performed best, but new capable models emerge frequently, reducing this dependency.
Construct validity: Section IV-A discussed the threat of ground-truth mapping mismatch. We mitigate by (i) using both ranked and unranked metrics, (ii) analyzing coverage per technique group, and (iii) qualitatively inspecting disagreements to separate genuine errors from plausible alternatives.
Conclusion validity: Att&ck evolves, and deprecated or merged techniques can yield ambiguity in labeling. Expert mapping at sub-technique levels is also more subjective. To mitigate, we lift sub-techniques to their top-level technique.
V-H Ethical considerations
We acknowledge that Triage has a measurable dual-use risk, considering that adversaries could use the predictions to prioritize exploitation paths or craft more effective payloads. Such dual-use risks are common for security-related tooling, e.g., a malware detector can be misused to assess sufficient obfuscation of malicious code. In the case of Triage, the risk is offset by substantial defensive benefits: by predicting the most likely Att&ck techniques for a vulnerability, it helps defenders prioritize patches, focus threat-hunting efforts, and map mitigations to realistic attack paths. As such, we believe that the benefits of providing defenders with capable tooling far outweigh the risks of misuse by attackers.
VI Concluding Remarks
This paper presented a variety of ways LLMs can be prompted to map CVEs to Att&ck techniques. We conclude that a hybrid combination of prompting with expert rules derived from the CMM and in-context learning yields the best performance, with in-context learning contributing most to the overall results. Our proposed approach, Triage, can automatically map a vulnerability to a list of potential Att&ck techniques, ranked by their likelihood, and it outperforms the current state-of-the-art in automated, ranked, mapping of CVEs.
Our ablation studies show that the number of in-context examples plays an important role in the performance. Moreover, our qualitative analysis reveals that, even when a CVE description explicitly mentions a vulnerability type associated with known attack techniques, the annotators may assign different techniques – or none at all. However, Triage was still able to predict plausible primary and secondary impacts (cf. the CMM), showing the benefit of the hybrid approach.
We do see that the approach underperforms in predicting secondary impacts, which we believe is in part due to not having enough labeled training samples of this type, but more importantly, because the CVE descriptions lack details that allow reasoning this deep into the CMM intrusion kill chain. This highlights the complexity of the mapping task and the limitations of relying solely on surface-level cues.
Future Work: For future work, we see a clear need for further standardization of datasets and annotation guidelines. In addition, we are interested in comparing our work to VTT-LLM [15]. A direct comparison is not possible, but one way to proceed is to make our approach uncategorized (like in Section IV-E) and evaluate it unranked. In addition, the datasets used by both approaches would need to be aligned. Finally, we would like to explore the use of reasoning models in Triage, examine to what extent LLMs can help analyze threat reports, and investigate using knowledge graphs to structure and integrate existing knowledge into this process.
Data Availability
To support open science and allow for replication and verification of our work, a replication package with prompts, scripts, and results is made available via Zenodo [6].
Acknowledgements
This work is supported by the Research Council of Norway through the secureIT project (#288787). The empirical evaluation made use of the Experimental Infrastructure for Exploration of Exascale Computing (eX3), financially supported by the Research Council of Norway under contract #270053.